### v2

# Core equality of real sequences

Abstract.
Given an ideal $\mathcal{I}$ on $\omega$ and a bounded real sequence $\bm{x}$, we denote by $\mathrm{core}_{\bm{x}}(\mathcal{I})$ the smallest interval $[a,b]$ such that $\{n\in\omega:x_{n}\notin[a-\varepsilon,b+\varepsilon]\}\in\mathcal{I}$ for all $\varepsilon>0$ (which corresponds to the interval [lim $[\,\liminf\bm{x},\limsup\bm{x}\,]$ if $\mathcal{I}$ is the ideal $\mathrm{Fin}$ of finite subsets of $\omega$).
First, we characterize all the infinite real matrices $A$ such that
$\mathrm{core}_{A\bm{x}}(\mathcal{J})=\mathrm{core}_{\bm{x}}(\mathcal{I})$
for all bounded sequences $\bm{x}$, provided that $\mathcal{J}$ is a countably generated ideal on $\omega$ and $A$ maps bounded sequences into bounded sequences. Such characterization fails if both $\mathcal{I}$ and $\mathcal{J}$ are the ideal of asymptotic density zero sets. Next, we show that such equality is possible for distinct ideals $\mathcal{I},\mathcal{J}$, answering an open question in [J. Math. Anal. Appl. 321 (2006), 515–523]. Lastly, we prove that, if $\mathcal{J}=\mathrm{Fin}$, the above equality holds for some matrix $A$ if and only if $\mathcal{I}=\mathrm{Fin}$ or $\mathcal{I}$ is an isomorphic copy of $\mathrm{Fin}\oplus\mathcal{P}(\omega)$ on $\omega$.
Key words and phrases:
Ideal convergence; summability; regular matrices; Rudin–Keisler order; ideal core.
2020 Mathematics Subject Classification:
Primary: 40A35, 40G15. Secondary: 40H05, 54A20.
1. Introduction
Let $\mathcal{I}$ be an ideal on the natural numbers $\omega$, that is, a family of subsets of $\omega$ closed under subsets and finite unions. Unless otherwise stated, it is assumed that $\mathcal{I}$ contains the family $\mathrm{Fin}$ of finite sets and that $\omega\notin\mathcal{I}$. We denote by $\mathcal{I}^{+}:=\mathcal{P}(\omega)\setminus\mathcal{I}$ and :=$\mathcal{I}^{\star}:=\{S\subseteq\omega:\omega\setminus S\in\mathcal{I}\}$ the family of $\mathcal{I}$-positive sets and the dual filter of $\mathcal{I}$, respectively. Ideals are regarded as subsets of the Cantor space $\{0,1\}^{\omega}$, hence it is possible to speak about $F_{\sigma}$-ideals, analytic ideals, meager ideals, etc. An important example of an ideal is the family $\mathcal{Z}$ of sets $S\subseteq\omega$ with asymptotic density zero, that is, $|S\cap[0,n]|=o(n)$ as $n\to\infty$.
Given a sequence $\bm{x}=(x_{n}:n\in\omega)$ taking values in a topological space $X$, we define its $\bm{\mathcal{I}}$-core by
$\mathrm{core}_{\bm{x}}(\mathcal{I}):=\bigcap_{S\in\mathcal{I}^{\star}}\overline{\mathrm{co}}\left\{x_{n}:n\in S\right\},$
where $\mathrm{co}$ denotes the convex hull operator and $\overline{\mathrm{co}}$ its closure, see [14, 22]. In the cases where $\mathcal{I}=\mathrm{Fin}$ and $\mathcal{I}=\mathcal{Z}$, the $\mathcal{I}$-core of $\bm{x}$ is usually called “Knopp core”  and “statistical core,”  respectively, see [12, 13, 18, 29]. Let also $\Gamma_{\bm{x}}(\mathcal{I})$ denote the set of $\mathcal{I}$-cluster points of $\bm{x}$, that is, the set of all $\eta\in X$ such that $\{n\in\omega:x_{n}\in U\}\in\mathcal{I}^{+}$ for all neighborhoods $U$ of $\eta$. It has been shown in [22, Corollary 2.3] that, if $\bm{x}$ is bounded real sequence, then $\mathrm{core}_{\bm{x}}(\mathcal{I})=\mathrm{co}(\Gamma_{\bm{x}}(\mathcal{I}))$. In addition, since $\Gamma_{\bm{x}}(\mathcal{I})$ is a nonempty compact set, see e.g. [25, Lemma 3.1], it follows that
$\mathrm{core}_{\bm{x}}(\mathcal{I})=\left[\,\mathcal{I}\text{-}\liminf\bm{x},\mathcal{I}\text{-}\limsup\bm{x}\,\right]$ (1)
[…]
Given an infinite matrix $A=(a_{n,k}:n,k\in\omega)$, we denote by $\mathrm{dom}(A)$ its domain, that is, the set of real sequences $\bm{x}=(x_{n})$ for which the $A$-transformed sequences $A\bm{x}:=\left(A_{n}\bm{x}:n\in\omega\right)$ are well defined, which means that the series
$A_{n}\bm{x}:=\sum_{k\in\omega}a_{n,k}x_{k}$
is convergent to a (finite) real number for each $n\in\omega$. We write $A\geq 0$ if $a_{n,k}\geq 0$ for all $n,k\in\omega$. Given sequence spaces $\mathcal{A},\mathcal{B}\subseteq\mathbf{R}^{\omega}$, we denote by $(\mathcal{A},\mathcal{B})$ the family of infinite matrices $A$ such that $\mathcal{A}\subseteq\mathrm{dom}(A)$ and $A\bm{x}\in\mathcal{B}$ for all $\bm{x}\in\mathcal{A}$. For instance, it is well known that $A\in(\ell_{\infty},\ell_{\infty})$ if and only if $A\in(c,c)$ if and only if $\|A\|<\infty$, where $\|A\|:=\sup_{n}\sum_{k}|a_{n,k}|$, see e.g. [5, Theorem 2.3.5]. Here, as usual, $\ell_{\infty}$ and $c$ stand for the vector space of bounded real sequences and convergent real sequences, respectively. $\ell_{\infty}$ and all its subspaces are endowed with the topology induced by the supremum norm.
[…]
$A\in(c_{b}(\mathcal{I}),c_{b}(\mathcal{J}))\quad\text{ and }\quad\mathcal{I}\text{-}\lim\bm{x}=\mathcal{J}\text{-}\lim A\bm{x}\,\,\text{ for all }\bm{x}\in c_{b}(\mathcal{I}),$
see e.g. [7, 19, 23]. Note that $(\mathrm{Fin},\mathrm{Fin})$-regular matrices are simply the classical regular matrices. Probably the most important regular matrix is the Cesàro matrix $C_{1}=(a_{n,k})$ defined by $(a_{n,k})=1/n$ if $k\leq n$ and $a_{n,k}=0$ otherwise.
A characterization of $(\mathcal{I},\mathcal{J})$-regular matrices has been recently proved by the author and Jeff Connor in [7, Theorem 1.2 and Theorem 1.3], see also [19, Corollary 2.11].
Theorem 1.1.
Let $A$ be an infinite real matrix and fix ideals $\mathcal{I}$, $\mathcal{J}$ on $\omega$. Suppose also that $A\geq 0$ or $\mathcal{I}=\mathrm{Fin}$ or $\mathcal{J}$ is countably generated.
Then $A$ is $(\mathcal{I},\mathcal{J})$-regular if and only if:
  1. (T1)
    $\|A\|<\infty$;
  2. (T2)
    $\mathcal{J}\text{-}\lim_{n}\sum_{k}a_{n,k}=1$;
  3. (T3)
    $\mathcal{J}\text{-}\lim_{n}\sum_{k\in E}|a_{n,k}|=0$ for all $E\in\mathcal{I}$.
In the statement above, recall that an ideal $\mathcal{J}$ on $\omega$ is countably generated if there exists a sequence $(Q_{k}:k\in\omega)$ of subsets of $\omega$ such that $S\in\mathcal{I}$ if and only if $S\subseteq\bigcup_{k\in F}Q_{k}$ for some finite $F\in\mathrm{Fin}$. Examples of countably generated ideals include $\mathrm{Fin}$ and the isomorphic copies on $\omega$ of $\mathrm{Fin}\times\{\emptyset\}:=\{S\subseteq\omega^{2}:\exists n\in\omega,S\subseteq[0,n]\times\omega\}$ and $\mathrm{Fin}\oplus\mathcal{P}(\omega):=\{S\subseteq\{0,1\}\times\omega:|S\cap(\{0\}\times\omega)|<\infty\}$, cf. [19, Remark 2.16]. Hereafter, an ideal $\mathcal{I}$ on a countably infinite set $W$ is said to be an isomorphic copy of an ideal $\mathcal{J}$ on $\omega$ if there exists a bijection $f:W\to\omega$ such that $S\in\mathcal{J}$ if and only if $f^{-1}[S]\in\mathcal{I}$ for each $S\subseteq\omega$.)
The above result extends the classical Silverman–Toeplitz characterization, which corresponds to the case $\mathcal{I}=\mathcal{J}=\mathrm{Fin}$. Lastly, it is worth mentioning that Theorem 1.1 does not hold for arbitrarily ideals: indeed, there exists a $(\mathcal{Z},\mathcal{Z})$-regular which does not satisfy condition (T3), see [7, Theorem 1.4].
2. Main results
Given ideals $\mathcal{I},\mathcal{J}$ on $\omega$, we study the core equality problem
$\mathrm{core}_{A\bm{x}}(\mathcal{J})=\mathrm{core}_{\bm{x}}(\mathcal{I})\quad\text{ for all sequences }\bm{x}\in\ell_{\infty}.$ (2)
More explicitly, we obtain necessary and sufficient conditions on the entries of $A$ to satisfy equality (2) and, then, we study the existence of such matrices $A$.
The first result in this direction has been obtained by Allen [1], which provides a characterization of the matrices $A$ which satisfy equality (2) in the case $\mathcal{I}=\mathcal{J}=\mathrm{Fin}$:
Theorem 2.1.
Let $A$ be an infinite real matrix and suppose that $\mathcal{I}=\mathcal{J}=\mathrm{Fin}$.
Then equality (2) holds if and only if:
  1. (A1)
    $A$ is regular;
  2. (A2)
    $\lim_{n}\sum_{k}|a_{n,k}|=1$;
  3. (A3)
    for each infinite $E\subseteq\omega$, there exists a strictly increasing sequence $(n_{i}:i\in\omega)$ in $\omega$ such that $\lim_{i}\sum_{k\in E}a_{n_{i},k}=1$
Note that, taking into account (A2), condition (A3) can be rewritten equivalently as $\limsup_{n}\sum_{k\in E}|a_{n,k}|=1$ for all infinite $E\subseteq\omega$. In addition, condition (A2) implies that $A$ maps bounded sequences into bounded sequences, i.e., $A\in(\ell_{\infty},\ell_{\infty})$.
A partial extension of the result above has been obtained by Connor, Fridy, and Orhan in the case where all the entries of $A$ are nonnegative, see [6, Theorem 2.1].
Theorem 2.2.
Let $A\in(\ell_{\infty},\ell_{\infty})$ be an infinite real matrix, let $\mathcal{I}$, $\mathcal{J}$ be ideals on $\omega$, and suppose that $A\geq 0$.
Then equality (2) holds if and only if:
  1. (C1)
    $A$ is $(\mathcal{I},\mathcal{J})$-regular;
  2. (C2)
    $\mathcal{J}\text{-}\limsup_{n}\sum_{k\in E}a_{n,k}=1$ for all $E\in\mathcal{I}^{+}$.
Our first main result removes the hypotheses that the entries of $A$ are nonnegative and, on the other hand, it requires that $\mathcal{J}$ is countably generated. Hence it provides a generalization of Theorem 2.1.
Theorem 2.3.
Let $A\in(\ell_{\infty},\ell_{\infty})$ be an infinite real matrix, let $\mathcal{I}$, $\mathcal{J}$ be ideals on $\omega$, and suppose that $\mathcal{J}$ is countably generated.
Then equality (2) holds if and only if:
  1. (L1)
    $A$ is $(\mathcal{I},\mathcal{J})$-regular;
  2. (L2)
    $\mathcal{J}\text{-}\limsup_{n}\sum_{k\in E}|a_{n,k}|=1$ for all $E\in\mathcal{I}^{+}$.
The proof of Theorem 2.3 recovers also Theorem 2.2, see Remark 3.2 below. In addition, the above characterization does not hold without any constraint on the ideals $\mathcal{I},\mathcal{J}$. Indeed, it fails for $\mathcal{I}=\mathcal{J}=\mathcal{Z}$, see Remark 3.3.
At this point, another result by Connor, Fridy, and Orhan proves that there are no regular matrices $A$ satisfying equality (2) if $\mathcal{I}=\mathcal{Z}$ and $\mathcal{J}=\mathrm{Fin}$, see [6, Theorem 2.4]. Accordingly, the authors left as open question whether there exist distinct ideals $\mathcal{I},\mathcal{J}$ on $\omega$ for which equality (2) holds for some matrix $A\in(\ell_{\infty},\ell_{\infty})$. Next, we show that the answer is affirmative. To this aim, recall that an ideal $\mathcal{I}$ is Rudin–Keisler below an ideal $\mathcal{J}$, shortened as $\mathcal{I}\leq_{\mathrm{RK}}\mathcal{J}$, if there exists a map $h:\omega\to\omega$ such that $S\in\mathcal{I}$ if and only if $h^{-1}[S]\in\mathcal{J}$.
Theorem 2.4.
Let $\mathcal{I}$, $\mathcal{J}$ be ideals on $\omega$ such that $\mathcal{I}\leq_{\mathrm{RK}}\mathcal{J}$. Then there exists an infinite real matrix $A\in(\ell_{\infty},\ell_{\infty})$ which satisfies equality (2).
It is worth to recall that, if $\mathcal{J}$ is a $P$-ideal (that is, for all increasing sequences $(J_{n})$ in $\mathcal{J}$ there exists $J\in\mathcal{J}$ such that $J_{n}\setminus J\in\mathrm{Fin}$ for all $n\in\omega$) and $\mathcal{I}\leq_{\mathrm{RK}}\mathcal{J}$, then it is possible to choose a witnessing function $h$ which is finite-to-one, see [9, Proposition 1.3.1]. Rudin–Keisler ordering and the latter stronger variant (known as Rudin–Blass ordering) on the maximal ideals are extensively studied in the literature, cf. [9, Section 1.3] and references therein. Examples of (distinct) ideals $\mathcal{I},\mathcal{J}$ on $\omega$ such that $\mathcal{I}\leq_{\mathrm{RK}}\mathcal{J}$ are abundant. For instance, it is known that $\mathrm{Fin}\leq_{\mathrm{RK}}\mathcal{J}$ for all meager ideals $\mathcal{J}$, see [9, Corollary 3.10.2]. In addition, $\mathcal{I}\leq_{\mathrm{RK}}\mathcal{J}$ for all Erdős–Ulam ideals $\mathcal{I},\mathcal{J}$ (where an ideal $\mathcal{H}$ is said to be Erdős–Ulam if there exists a nonnegative real sequence $(h_{n})$ such that $\sum_{n}h_{n}=\infty$, $h_{n}=o(\sum_{k\leq n}h_{k})$ as $n\to\infty$, and $S\in\mathcal{H}$ if and only if $\sum_{k\in S,k\leq n}h_{k}=o(\sum_{k\leq n}h_{k})$ as $n\to\infty$), see [9, Lemma 1.13.10], cf. also [17, Corollary 1]. Other examples can be found within the class of summable ideals, see [9, Section 1.12].
Our last result extends the latter [6, Theorem 2.4] by finding all ideals $\mathcal{I}$ for which equality (2) holds with $\mathcal{J}=\mathrm{Fin}$ and some matrix $A$:
Theorem 2.5.
There exists an infinite real matrix $A\in(\ell_{\infty},\ell_{\infty})$ satisfying equality (2) with $\mathcal{J}=\mathrm{Fin}$ if and only if $\mathcal{I}=\mathrm{Fin}$ or $\mathcal{I}$ is an isomorphic copy of $\mathrm{Fin}\oplus\mathcal{P}(\omega)$ on $\omega$.
(Equivalently, the latter condition means =$\mathcal{I}=\{S\subseteq\omega:S\cap T\in\mathrm{Fin}\}$ for some infinite $T\subseteq\omega$.) It is worth noting that Theorem 2.5 is also related to the question posed by Mazur in The Scottish Book whether the notion of statistical convergence (i.e., $\mathcal{Z}$-convergence) of bounded sequences is equivalent to some matrix summability method, see [10] and references therein. A positive answer has been given by Khan and Orhan in [16, Theorem 2.2].
Based on the previous observations, we leave as an open question to check whether there exists a matrix $A\in(\ell_{\infty},\ell_{\infty})$ satisfying equality (2) with $\mathcal{I}=\mathrm{Fin}$ if and only if $\mathcal{J}$ is meager. In the same direction, it would be interesting to know if the condition $\mathcal{I}\leq_{\mathrm{RK}}\mathcal{J}$ is also necessary in the statament of Theorem 2.4.
3. Proofs
We start with an auxiliary lemma and, then, we proceed to the proofs of our results.
Lemma 3.1.
Let $\bm{x}$ and $\bm{y}$ be two relatively compact sequences taking values in a locally convex topological vector space $X$. Let $\mathcal{I}$ be an ideal on $\omega$. Then
$\mathrm{core}_{\bm{x}}(\mathcal{I})=\mathrm{core}_{\bm{y}}(\mathcal{I})$
whenever -$\mathcal{I}\text{-}\lim(\bm{x}-\bm{y})=0$.
Proof.
Since $X$ is, in particular, a topological group, it follows by [25, Lemma 3.5] and the hypothesis -$\mathcal{I}\text{-}\lim(\bm{x}-\bm{y})=0$ that $\Gamma_{\bm{x}}(\mathcal{I})=\Gamma_{\bm{y}}(\mathcal{I})$. Now, let $K_{\bm{x}}$ be the closure of the image $\{x_{n}:n\in\omega\}$, which is a compact subset of the locally convex space $X$. Of course, this implies that $\{x_{n}:n\in\omega\}\subseteq K_{\bm{x}}+U$ for every open neighborhood $U$ of 00 (which means that $\bm{x}$ is “$\mathcal{I}^{\star}$-asymptotically $K_{\bm{x}}$-controlled,” using the terminology in [14, Definition 3.2]). It follows by [14, Theorem 3.4] that $\mathrm{core}_{\bm{x}}(\mathcal{I})=\overline{\mathrm{co}}\,\Gamma_{\bm{x}}(\mathcal{I})$; cf. also [22, Theorem 2.2] in the case of first countable locally convex spaces. With an analog reasoning on the sequence $\bm{y}$, we conclude that $\mathrm{core}_{\bm{x}}(\mathcal{I})=\overline{\mathrm{co}}\,\Gamma_{\bm{x}}(\mathcal{I})=\overline{\mathrm{co}}\,\Gamma_{\bm{y}}(\mathcal{I})=\mathrm{core}_{\bm{y}}(\mathcal{I})$. ∎
Proof of Theorem 2.3.
Only If part. Pick a sequence $\bm{x}\in c_{b}(\mathcal{I})$ and define $\eta:=\mathcal{I}\text{-}\lim\bm{x}$, so that $\Gamma_{\bm{x}}(\mathcal{I})=\{\eta\}$. It follows that $A\bm{x}$ is well-defined bounded sequence and, thanks to equality (2), that $\mathrm{core}_{A\bm{x}}(\mathcal{J})=\mathrm{core}_{\bm{x}}(\mathcal{I})=\{\eta\}$. We conclude by [22, Proposition 4.2] that $\mathcal{J}\text{-}\lim A\bm{x}=\eta$, therefore $A$ is $(\mathcal{I},\mathcal{J})$-regular.
At this point, pick a set $E\in\mathcal{I}^{+}$. Recall that the hypothesis $A\in(\ell_{\infty},\ell_{\infty})$ is equivalent to $\sup_{n\in\omega}\sum_{k}|a_{n,k}|<\infty$, see e.g. [5, Theorem 2.3.5]. In addition, it follows by Theorem 1.1 that $\mathcal{J}\text{-}\lim_{n}|a_{n,k}|=0$ for all $k\in\omega$. This implies that conditions (K1)–(K3) in the statement of [19, Corollary 4.3] are satisfied (in the case where $d=m=1$, so that $a_{n,k}(i,j)$ is simply $a_{n,k}$ considering that both $i$ and $j$ can take only one value). Hence, thanks to [19, Corollary 4.3], there exists a $\{-1,0,1\}$-valued sequence $\bm{x}$ supported on $E$ such that
$\mathcal{J}\text{-}\limsup_{n\to\infty}\sum_{k\in E}|a_{n,k}|=\mathcal{J}\text{-}\limsup_{n\to\infty}\left|A_{n}\bm{x}\right|.$ (3)
(Again, equation (3) is a rewriting of the claim in [19, Corollary 4.3] taking into account that both variables $i$ and $j$ can take only value.) Define $F:=\{n\in E:x_{n}=1\}$ and $G:=\{n\in E:x_{n}=-1\}$, so that $\{F,G\}$ is a partition of $E$ and $\bm{x}:=\bm{1}_{F}-\bm{1}_{G}$ (hereafter, $\bm{1}_{S}$ stands for the characteristic function of $S$). Note that, since $\mathcal{I}$ is an ideal, then at least one between $F$ and $G$ is an $\mathcal{I}$-positive set. For convenience, let $h:\mathbf{R}\to\mathbf{R}$ be the function defined by $\eta\mapsto|\eta|$. Since $h$ is continuous, it follows by [24, Proposition 3.2] that
$\Gamma_{h(A\bm{x})}(\mathcal{J})=h\left[\Gamma_{A\bm{x}}(\mathcal{J})\right],$
[…]
$\begin{split}\mathcal{J}\text{-}\limsup_{n\to\infty}\sum_{k\in E}|a_{n,k}|&=\max\,\Gamma_{h(A\bm{x})}(\mathcal{J})=\max\,h\left[\Gamma_{A\bm{x}}(\mathcal{J})\right]\\ &=\max\{|\mathcal{J}\text{-}\limsup A\bm{x}|,|\mathcal{J}\text{-}\liminf A\bm{x}|\}\\ &=\max\{|\mathcal{J}\text{-}\limsup\bm{x}|,|\mathcal{J}\text{-}\liminf\bm{x}|\}=\max\,h\left[\Gamma_{\bm{x}}(\mathcal{J})\right]=1.\end{split}$
Therefore both conditions (L1) and (L2) hold.
If part. Conversely, let $A=(a_{n,k}:n,k\in\omega)\in(\ell_{\infty},\ell_{\infty})$ be a $(\mathcal{I},\mathcal{J})$-regular matrix which satisfies condition (L2). Then we get
$1=\mathcal{J}\text{-}\lim_{n\to\infty}\sum_{k\in\omega}a_{n,k}\leq\mathcal{J}\text{-}\liminf_{n\to\infty}\sum_{k\in\omega}|a_{n,k}|\leq\mathcal{J}\text{-}\limsup_{n\to\infty}\sum_{k\in\omega}|a_{n,k}|=1,$
so that $\mathcal{J}\text{-}\lim_{n}\sum_{k}|a_{n,k}|=\mathcal{J}\text{-}\lim_{n}\sum_{k}a_{n,k}=1$. Decomposing each $a_{n,k}$ into its positive and negative part as $a_{n,k}^{+}-a_{n,k}^{-}$ for all $n,k\in\omega$, it follows that
$\mathcal{J}\text{-}\lim_{n\to\infty}\sum_{k\in\omega}a_{n,k}^{-}=0\quad\text{ and }\quad\mathcal{J}\text{-}\lim_{n\to\infty}\sum_{k\in\omega}a_{n,k}^{+}=1.$ (4)
At this point, pick a sequence $\bm{x}\in\ell_{\infty}$, define $\eta:=\mathcal{I}\text{-}\limsup\bm{x}$. Considering that $\mathrm{core}_{\bm{x}+\kappa\bm{1}_{\omega}}(\mathcal{I})=\mathrm{core}_{\bm{x}}(\mathcal{I})+\{\kappa\}$ and also, by the $(\mathcal{I},\mathcal{J})$-regularity of $A$, that $\mathrm{core}_{A(\bm{x}+\kappa\bm{1}_{\omega})}(\mathcal{J})=\mathrm{core}_{A\bm{x}}(\mathcal{J})+\{\kappa\}$ for all $\kappa\in\mathbf{R}$, we can suppose without loss of generality that $\eta>0$. Fix an arbitrary $\varepsilon>0$ and define
$\delta:=\min\left\{\,\frac{\varepsilon}{2+\eta+4\|x\|},\,1\right\}\quad\text{ and }\quad E:=\left\{k\in\omega:x_{k}\geq\eta-\delta\right\}.$
Note that $\delta>0$ and $E$ is an $\mathcal{I}$-positive set since $\eta$ is an $\mathcal{I}$-cluster point of $\bm{x}$. It follows by (4) and condition (L2) that $\mathcal{J}\text{-}\limsup_{n}\sum_{k\in E}a_{n,k}^{+}=1$. Thus, define
$S:=\left\{n\in\omega:1-\delta\leq\sum_{k\in E}a_{n,k}^{+}\leq\sum_{k\in\omega}|a_{n,k}|\leq 1+\delta\right\}.$ (5)
Observe the first inequality in the definition of $S$ holds on a $\mathcal{J}$-positive set, the second one for all $n$, and the latter one on $\mathcal{J}^{\star}$. Therefore $S\in\mathcal{J}^{+}$. For each $n\in S$, it also follows that $|\sum_{k\in\omega}a_{n,k}^{+}-1|\leq\delta$, so that
$\left|\,\sum_{k\in\omega}a_{n,k}^{-}\,\right|\leq 2\delta\quad\text{ and }\quad\left|\,\sum_{k\in E^{c}}a_{n,k}^{+}\,\right|\leq 2\delta.$ (6)
Putting all together, we obtain that, for all $n\in S$,
$\begin{split}A_{n}\bm{x}&=\sum_{k\in E}a_{n,k}^{+}x_{k}+\sum_{k\in E^{c}}a_{n,k}^{+}x_{k}-\sum_{k\in\omega}a_{n,k}^{-}x_{k}\\ &\geq(1-\delta)(\eta-\delta)-2\delta\|x\|-2\delta\|x\|\\ &\geq\eta-\delta(1+\eta+4\|x\|)\geq\eta-\varepsilon.\end{split}$
At the same time, define $E^{\prime}:=\{k\in\omega:x_{k}\leq\eta+\delta\}$, which belongs to $\mathcal{I}^{\star}$ and note, similarly as above, that $\mathcal{J}\text{-}\lim_{n}\sum_{k\in E^{\prime}}|a_{n,k}|=\mathcal{J}\text{-}\lim_{n}\sum_{k\in E^{\prime}}a_{n,k}=\mathcal{J}\text{-}\lim_{n}\sum_{k\in E^{\prime}}a_{n,k}^{+}=1$ and $\mathcal{J}\text{-}\lim_{n}\sum_{k\in\omega}a_{n,k}^{-}=0$. Let $S^{\prime}$ be the set defined as in (5) replacing $E$ with $E^{\prime}$, and note that $S^{\prime}\in\mathcal{J}^{\star}$. Similarly, estimates (6) hold for all $n\in S^{\prime}$ replacing $E$ with $E^{\prime}$. Putting again all together, we obtain that, for all $n\in S^{\prime}$,
$\begin{split}A_{n}\bm{x}&=\sum_{k\in E}a_{n,k}^{+}x_{k}+\sum_{k\in E^{c}}a_{n,k}^{+}x_{k}-\sum_{k\in\omega}a_{n,k}^{-}x_{k}\\ &\leq(1+\delta)(\eta+\delta)+2\delta\|x\|+2\delta\|x\|\\ &\leq\eta+\delta(2+\eta+4\|x\|)\leq\eta+\varepsilon.\end{split}$
[…]
Remark 3.2.
It is clear from the proof above that the If part holds without any additional hypothesis on $\mathcal{J}$. Moreover, the fact the $\mathcal{J}$ is countably generated has been used only once in the proof of the Only If part, precisely in the existence of a $\{-1,0,1\}$-valued sequence $\bm{x}$ supported on a given $E\in\mathcal{I}^{+}$ and satisfying equality (3). The latter is trivial if $A\geq 0$ by choosing $\bm{x}=\bm{1}_{E}$. In this sense, we recover also Theorem 2.2.
Remark 3.3.
On the other hand, if $\mathcal{I}=\mathcal{J}=\mathcal{Z}$, then the analogue of Theorem 2.3 does not hold. Indeed, thanks to the proof of [7, Theorem 1.4] there exists a matrix $A\in(\ell_{\infty},c_{0}(\mathcal{Z})\cap\ell_{\infty})$ and an infinite set $I\in\mathcal{I}$ such that
$\mathcal{Z}\text{-}\limsup_{n\to\infty}\sum_{k\in I}|a_{n,k}|=1.$
(Here, $c_{0}(\mathcal{Z})$ stands for the vector space of sequences which are $\mathcal{Z}$-convergent to 00.) At this point, define $B:=A+\mathrm{Id}$, where $\mathrm{Id}$ stands for the infinite identity matrix. On one hand, for each $\bm{x}\in\ell_{\infty}$ we have $B\bm{x}=A\bm{x}+\bm{x}$ and $A\bm{x}\in c_{0}(\mathcal{Z})\cap\ell_{\infty}$, hence by Lemma 3.1 we get $\mathrm{core}_{B\bm{x}}(\mathcal{Z})=\mathrm{core}_{\bm{x}}(\mathcal{Z})$. Thus equality (2) holds for the matrix $B$. On the other hand,
$\mathcal{Z}\text{-}\limsup_{n\to\infty}\sum_{k\in\omega}|b_{n,k}|=1+\mathcal{Z}\text{-}\limsup_{n\to\infty}\sum_{k\in\omega}|a_{n,k}|\geq 1+\mathcal{Z}\text{-}\limsup_{n\to\infty}\sum_{k\in I}|a_{n,k}|=2.$
This shows that $B$ does not satisfy condition (L2).
Proof of Theorem 2.4.
Since $\mathcal{I}\leq_{\mathrm{RK}}\mathcal{J}$, there exists a map $h:\omega\to\omega$ such that $S\in\mathcal{I}$ if and only if $h^{-1}[S]\in\mathcal{J}$. Now, let $A=(a_{n,k}:n,k\in\omega)$ be the matrix defined by
$a_{n,k}=\begin{cases}\,1\,\,\,\,&\text{ if }k=h(n),\\ \,0&\text{ otherwise}.\end{cases}$
Note that $A\in(\ell_{\infty},\ell_{\infty})$ since every row contains a single $1$ (however, $A$ is not necessarily regular if the witnessing map $h$ cannot be chosen finite-to-one). Fix $\bm{x}\in\ell_{\infty}$, let $U\subseteq\mathbf{R}$ be a nonempty open set, and define $S:=\{n\in\omega:x_{n}\in U\}$. Observe also that $A_{n}\bm{x}=\sum_{k}a_{n,k}x_{k}=x_{h(n)}$ for all $n\in\omega$. It follows that $S\in\mathcal{I}$ if and only if
$h^{-1}[S]=\{n\in\omega:x_{h(n)}\in U\}=\{n\in\omega:A_{n}\bm{x}\in U\}\in\mathcal{J}.$
This implies that $\Gamma_{\bm{x}}(\mathcal{I})=\Gamma_{A\bm{x}}(\mathcal{J})$, so that by [22, Corollary 2.3] we get
$\mathrm{core}_{\bm{x}}(\mathcal{I})=\mathrm{co}(\Gamma_{\bm{x}}(\mathcal{I}))=\mathrm{co}(\Gamma_{A\bm{x}}(\mathcal{J}))=\mathrm{core}_{A\bm{x}}(\mathcal{J}).$
Therefore equality (2) holds. ∎
For our last proof, we need to recall that an ideal $\mathcal{I}$ on $\omega$ is said to be:
  1. (i)
    a $P$-ideal if for all increasing sequences $(I_{n})$ in $\mathcal{I}$ there exists $I\in\mathcal{I}$ such that $I_{n}\setminus I\in\mathrm{Fin}$ for all $n\in\omega$;
  2. (ii)
    a $P^{+}$-ideal if for all decreasing sequences $(I_{n})$ in $\mathcal{I}^{+}$ there exists $I\in\mathcal{I}^{+}$ such that $I\setminus I_{n}\in\mathrm{Fin}$ for all $n\in\omega$;
  3. (iii)
    tall if every infinite set $S\subseteq\omega$ contains an infinite subset $I\subseteq S$ such that $I\in\mathcal{I}$;
  4. (iv)
    nowhere tall if, for every $S\in\mathcal{I}^{+}$, the ideal $\mathcal{I}\upharpoonright S:=\mathcal{I}\cap\mathcal{P}(S)$ is not tall.
Proof of Theorem 2.5.
If part. Suppose that =$\mathcal{I}=\{S\subseteq\omega:S\cap T\in\mathrm{Fin}\}$ for some infinite $T\subseteq\omega$ (i.e., $\mathcal{I}=\mathrm{Fin}$ if $T$ is cofinite or $\mathcal{I}$ is an isomorphic copy of $\mathrm{Fin}\oplus\mathcal{P}(\omega)$ on $\omega$ otherwise). Then $\mathcal{I}\leq_{\mathrm{RK}}\mathrm{Fin}$. In fact, if $(t_{n}:n\in\omega)$ denotes the increasing enumeration of elements of $T$, one can choose the witnessing map $h:\omega\to\omega$ defined by $h(n)=t_{n}$ for all $n\in\omega$. The claim follows by Theorem 2.4.
Only If part. Let $\mathcal{I}$ be an ideal on $\omega$ for which there exists a matrix $A\in(\ell_{\infty},\ell_{\infty})$ which satisfies equality (2) with $\mathcal{J}=\mathrm{Fin}$. We divide the remaining proof in several claims.
Claim 1.
$\mathcal{I}$ is an analytic ideal.
Proof.
Observe that, for each $E\subseteq\omega$, we have $E\in\mathcal{I}$ if and only if -lim $\mathcal{I}\text{-}\limsup\bm{1}_{E}\leq 0$. It follows by equality (2) that
=$\mathcal{I}=\left\{E\subseteq\omega:\limsup A\bm{1}_{E}\leq 0\right\}=\bigcap_{p\in\omega}\bigcup_{q\in\omega}\bigcap_{n\geq q}G_{p,n},$
where $G_{p,n}:=\left\{E\subseteq\omega:\sum_{k\in E}a_{n,k}<2^{-p}\right\}$ for all $n,p\in\omega$. Hence, it is sufficient to show that each $G_{n,p}$ is open. For, fix $n,p\in\omega$. If $G_{n,p}=\emptyset$, then it is open. Otherwise fix $E\in G_{n,p}$. Since $\sum_{k}|a_{n,k}|\leq\|A\|<\infty$, there exists $k_{0}\in\omega$ such that
$\sum_{k>k_{0}}|a_{n,k}|<\frac{1}{2}\left(2^{-p}-\sum_{k\in E}a_{n,k}\right).$
Now, let $F\subseteq\omega$ be a set such that $E\cap[0,k_{0}]=F\cap[0,k_{0}]$. It follows that
$\begin{split}\sum_{k\in F}a_{n,k}&\leq\sum_{k\in F\cap[0,k_{0}]}a_{n,k}+\sum_{k\in F\setminus[0,k_{0}]}|a_{n,k}|\\ &\leq\sum_{k\in E\cap[0,k_{0}]}a_{n,k}+\sum_{k>k_{0}}|a_{n,k}|\\ &\leq\sum_{k\in E}a_{n,k}+2\sum_{k>k_{0}}|a_{n,k}|<2^{-p}.\\ \end{split}$
This shows that $F\in G_{n,p}$. Hence $E$ is an interior point, so that $G_{n,p}$ is open. Therefore $\mathcal{I}$ is a $G_{\delta\sigma\delta}$-ideal. ∎
Claim 2.
$\mathcal{I}$ is a $P^{+}$-ideal.
Proof.
Let us suppose for the sake of contradiction that $\mathcal{I}$ is not a $P^{+}$-ideal, hence it is possible to fix a strictly decreasing sequence $(I_{n}:n\in\omega)$ in $\mathcal{I}^{+}$ such that for all sequences $(F_{n}:n\in\omega)$ of finite sets with $F_{n}\subseteq I_{n}\setminus I_{n+1}$ we have $\bigcup_{n}F_{n}\in\mathcal{I}$. Since $I_{0}\in\mathcal{I}^{+}$ it follows by equality (2) that lim $\limsup A\bm{1}_{I_{0}}=1$. Hence there exists $n_{0}\in\omega$ such that $A_{n_{0}}\bm{1}_{I_{0}}>1-2^{-0}$. Set $p_{0}:=0$ and pick an integer $q_{0}>p_{0}$ such that $\sum_{k\in F_{0}}a_{n_{0},k}\bm{1}_{I_{0}}(k)>1-2^{-0}$, where $F_{0}:=\omega\cap[p_{0},q_{0}]$. Recall also that $A$ is $(\mathcal{I},\mathrm{Fin})$-regular since it satisfies (2), hence it is regular. In particular, by Theorem 1.1, we have $\lim_{k}a_{n,k}=0$ and $\sum_{k}|a_{n,k}|<\infty$ for all $n,k\in\omega$. At this point, suppose that $n_{i-1}$ and $F_{i-1}:=\omega\cap[p_{i-1},q_{i-1}]$ have been defined for some positive integer $i$. Then, proceed recursively as follows:
  1. (i)
    Pick an integer $p_{i}>q_{i-1}$ with the property that
    $\sum_{k\geq p_{i}}|a_{n_{j},k}|<2^{-i}$
    for all $j\in\omega\cap[0,i-1]$.
  2. (ii)
    Let $n_{i}>n_{i-1}$ be an integer such that
    $A_{n_{i}}\bm{1}_{I_{i}}>1-2^{-i}\quad\text{ and }\quad\sum_{k<p_{i}}|a_{n,k}|<2^{-i}$
    for all integers $n\geq n_{i}$.
  3. (iii)
    Let $q_{i}>p_{i}$ be an integer such that
    $\sum_{k\in F_{i}}a_{n_{i},k}\bm{1}_{I_{i}}(k)>1-2^{1-i},$
    where $F_{i}:=\omega\cap[p_{i},q_{i}]$. (Note that this is possible because $\sum_{k\geq p_{i}}a_{n_{i},k}\bm{1}_{I_{i}}(k)$ is at least $A_{n_{i}}\bm{1}_{I_{i}}-\sum_{k<p_{i}}|a_{n_{i},k}|>1-2^{1-i}$.)
To conclude, define $F:=\bigcup_{i}F_{i}\cap I_{i}$. By the standing hypothesis, we have $F\in\mathcal{I}$, hence by equality (2) we get lim $\limsup A\bm{1}_{F}=\mathcal{I}\text{-}\limsup\bm{1}_{F}=0$. On the other hand, it follows by the construction above that, for all $i\geq 1$,
$\begin{split}\limsup_{n\to\infty}A_{n}\bm{1}_{F}&\geq\limsup_{i\to\infty}\sum_{k\in F}a_{n_{i},k}\\ &\geq\limsup_{i\to\infty}\left(\sum_{k\in F_{i}\cap I_{i}}a_{n_{i},k}-\sum_{k<p_{i}}|a_{n_{i},k}|-\sum_{k\geq p_{i+1}}|a_{n_{i},k}|\right)\\ &\geq\limsup_{i\to\infty}(1-2^{1-i}-2^{-i}-2^{-1-i})=1.\end{split}$
This contradiction proves that $\mathcal{I}$ is a $P^{+}$-ideal. ∎
Claim 3.
$\mathcal{I}$ is a $P$-ideal.
Proof.
Let us suppose for the sake of contradiction that $\mathcal{I}$ is not a $P$-ideal, hence it is possible to fix an increasing sequence $(I_{n}:n\in\omega)$ in $\mathcal{I}$ such that, for all sequences $(F_{n}:n\in\omega)$ with $F_{n}\subseteq D_{n}:=I_{n+1}\setminus I_{n}$ for each $n$, we have $\bigcup_{n}(D_{n}\setminus F_{n})\in\mathcal{I}^{+}$. Without loss of generality, we can assume that $I_{0}=\emptyset$. Define
$S:=\{n\in\omega:D_{n}\notin\mathrm{Fin}\}.$
It is easy to see that, if $S$ is finite, then $(I_{n}:n\in\omega)$ cannot be a sequence witnessing that $\mathcal{I}$ is not a $P$-ideal: indeed, in such case, $I:=I_{0}$ if $S=\emptyset$ or $I:=I_{1+\max S}$ if $S\neq\emptyset$ satisfies $I_{n}\setminus I\in\mathrm{Fin}$ for all $n\in\omega$. Hence $S$ has to be infinite, which implies that, passing to a suitable subsequence, we can assume without loss of generality that $D_{n}$ is infinite for all $n\in\omega$.
Now, note that, since each $I_{n}$ belongs to $\mathcal{I}$, then $\lim A\bm{1}_{I_{n}}=0$ for all $n\in\omega$ by equality (2). Let $(k_{n}:n\in\omega)$ be a strictly increasing sequence in $\omega$ such that
$\sum_{k>k_{n}}|a_{n,k}|<2^{-n}\quad\text{ for all }n\in\omega.$ (7)
It follows by the If part in the proof of Theorem 2.3 with $\mathcal{J}=\mathrm{Fin}$ that $\lim_{n}\sum_{k}a_{n,k}^{-}=0$. In particular, there exists a strictly increasing sequence $(h_{m}:m\in\omega)$ such that
$\sum_{k\in\omega}a_{n,k}^{-}<2^{-m}\quad\text{ for all }n\geq h_{m}.$ (8)
Let also $(t_{m}:m\in\omega)$ be a strictly increasing sequence in $\omega$ such that
$t_{m}\geq h_{m}\text{ and }\left|\,\sum_{i\leq m}A_{n}\bm{1}_{D_{i}}\right|<2^{-m}\quad\text{ for all }m\in\omega\text{ and }n\geq t_{m}.$ (9)
To conclude, define $F_{n}:=\omega\cap[0,k_{t_{n}}]$ for all $n\in\omega$ and set $D_{\infty}:=\bigcup_{n}(D_{n}\setminus F_{n})$. On the one hand, it follows by the standing hypothesis that $D_{\infty}\in\mathcal{I}^{+}$, hence by equality (2) we have lim $\limsup A\bm{1}_{D_{\infty}}=1$. On the other hand, pick $m\in\omega$ and fix $n\in[t_{m},t_{m+1})$. It follows that
$\begin{split}\left|A_{n}\bm{1}_{D_{\infty}}\right|&=\left|\,\sum_{i\in\omega}A_{n}\bm{1}_{D_{i}\setminus F_{i}}\right|\\ &\leq\sum_{k>k_{n}}|a_{n,k}|+\left|\,\sum_{i\in\omega}\sum_{k\leq k_{n}}a_{n,k}\bm{1}_{D_{i}\setminus F_{i}}(k)\,\right|\\ &\leq 2^{-n}+\left|\,\sum_{k\leq k_{n}}\sum_{i\in\omega}a_{n,k}\bm{1}_{D_{i}\setminus F_{i}}(k)\right|,\\ \end{split}$
where at the last inequality we used (7). At this point, notice that, if $k\leq k_{n}$ and $i>m$ then $\bm{1}_{D_{i}\setminus F_{i}}(k)=0$ since $\min(D_{i}\setminus F_{i})>\max(F_{i})=k_{t_{i}}\geq k_{t_{m+1}}>k_{n}$. Taking into account that $n\geq t_{m}\geq m$, inequality (8), and that $t_{m}\geq h_{m}$, we obtain
$\begin{split}\left|A_{n}\bm{1}_{D_{\infty}}\right|&\leq 2^{-m}+\left|\,\sum_{k\leq k_{n}}\sum_{i\leq m}a_{n,k}\bm{1}_{D_{i}\setminus F_{i}}(k)\right|\\ &\leq 2^{-m}+\left|\,\sum_{k\leq k_{n}}\sum_{i\leq m}a_{n,k}\bm{1}_{D_{i}}(k)\right|+\sum_{k\in\omega}a_{n,k}^{-}\\ &\leq 2^{1-m}+\left|\,\sum_{k\leq k_{n}}\sum_{i\leq m}a_{n,k}\bm{1}_{D_{i}}(k)\right|.\end{split}$
Lastly, using also inequality (9), we get
$\begin{split}\left|A_{n}\bm{1}_{D_{\infty}}\right|&\leq 2^{1-m}+\left|\,\sum_{k\in\omega}\sum_{i\leq m}a_{n,k}\bm{1}_{D_{i}}(k)\right|+\left|\sum_{k>k_{n}}\sum_{i\leq m}a_{n,k}\bm{1}_{D_{i}}(k)\right|\\ &\leq 2^{1-m}+\left|\,\sum_{i\leq m}A_{n}\bm{1}_{D_{i}}\right|+\sum_{k>k_{n}}|a_{n,k}|\\ &\leq 2^{1-m}+2^{-m}+2^{-n}\leq 4^{1-m}.\end{split}$
This proves that $\lim A\bm{1}_{D_{\infty}}=0$, which gives the desired contradiction. ∎
Thanks to Claims 1, 2, and 3, $\mathcal{I}$ is an analytic $P$-ideal which is also a $P^{+}$-ideal. Although it will not be used in the following results, it follows by [3, Theorem 2.5] that $\mathcal{I}$ is a necessarily $F_{\sigma}$ $P$-ideal (we omit details).
Claim 4.
$\mathcal{I}$ is not tall.
Proof.
Let us suppose for the sake of contradiction that $\mathcal{I}$ is tall. Define the infinite matrix $A^{+}:=(a_{n,k}^{+}:n,k\in\omega)$. Since $\lim_{n}\sum_{k}a_{n,k}^{-}=0$ (cf. the proof of Claim 3), it follows that $A^{+}$ is a nonnegative $(\mathcal{I},\mathrm{Fin})$-regular matrix; in particular, it is a nonnegative regular matrix. At this point, define the map $\mu^{\star}:\mathcal{P}(\omega)\to\mathbf{R}$ by
$\mu^{\star}(S):=\limsup A^{+}\bm{1}_{S}\quad\text{ for all }S\subseteq\omega.$
Note also that $\lim(A\bm{x}-A^{+}\bm{x})=0$ for all $\bm{x}\in\ell_{\infty}$, hence by Lemma 3.1 and equality (2)
$\mathrm{core}_{A^{+}\bm{x}}(\mathrm{Fin})=\mathrm{core}_{\bm{x}}(\mathcal{I})\quad\text{ for all sequences }\bm{x}\in\ell_{\infty}.$
Thus =$\mathcal{I}=\{S\subseteq\omega:\mathcal{I}\text{-}\lim\bm{1}_{S}=0\}=\{S\subseteq\omega:\lim A^{+}\bm{1}_{S}=0\}=\{S\subseteq\omega:\mu^{\star}(S)=0\}$. Since $\mathcal{I}$ is not tall, it follows by [8, Proposition 7.2] that $\lim_{n}\sup_{k}a_{n,k}^{+}=0$. In addition, recalling that $A^{+}$ is a nonnegative matrix, we have also that $\lim_{n}a_{n,k}^{+}=0$ and $\sum_{k}a_{n,k}^{+}<\infty$ for all $n,k\in\omega$ by Theorem 1.1. It follows by [8, Theorem 6.2] that the function $\mu^{\star}$ has the weak Darboux property, i.e., for each $S\subseteq\omega$ and $y\in[0,\mu^{\star}(S)]$ there exists $X\subseteq S$ such that $\mu^{\star}(X)=y$, cf. [27, Section 2]. This implies that there exists a decreasing sequence $(I_{m}:m\in\omega)$ of subsets of $\omega$ such that
$\mu^{\star}(I_{m})=2^{-m}\quad\text{ for all }m\in\omega.$
At this point, let $I\subseteq\omega$ be a set such that $J_{m}:=I\setminus I_{m}\in\mathrm{Fin}$ for all $m\in\omega$. Since $\mu^{\star}$ is monotone and subadditive, we obtain
$\mu^{\star}(I)\leq\mu^{\star}(I_{m})+\mu^{\star}(J_{m})=2^{-m}+\limsup_{n\to\infty}\sum_{k\in J_{m}}a_{n,k}=2^{-m}\quad\text{ for all }m\in\omega.$
Hence $\mu^{\star}(I)=0$, i.e., $I\in\mathcal{I}$. This proves that $\mathcal{I}$ is not a $P^{+}$-ideal, which contradicts Claim 2. Therefore $\mathcal{I}$ cannot be tall. ∎
Claim 5.
$\mathcal{I}$ is a nowhere tall ideal.
Proof.
Fix a set $S\in\mathcal{I}^{+}$ and consider the ideal ~$\tilde{\mathcal{I}}:=\mathcal{I}\upharpoonright S$. Since $\mathcal{I}$ is analytic by Claim 1 and $\mathcal{P}(S)$ is closed, then $\tilde{\mathcal{I}}$ is analytic as well. Moreover, since $\mathcal{I}$ is both a $P$-ideal and $P^{+}$-ideals by Claims 2 and 3, respectively, it is immediate that the same properties hold for $\tilde{\mathcal{I}}$. Let $\tilde{A}=(\tilde{a}_{n,k}:n,k\in\omega)$ be the matrix defined by $\tilde{a}_{n,k}:=a_{n,k}$ if $k\in S$ and $\tilde{a}_{n,k}:=0$ otherwise. Now, note that, by equality (2),
~=$\begin{split}\tilde{\mathcal{I}}&=\{X\subseteq S:\mathcal{I}\text{-}\lim\bm{1}_{X}=0\}=\{X\subseteq S:\lim A\bm{1}_{X}=0\}\\ &=\{X\subseteq S:\lim\tilde{A}\bm{1}_{X}=0\}=\{X\subseteq S:\tilde{\mu}^{\star}(X)=0\},\end{split}$
where $\tilde{\mu}^{\star}(X):=\limsup_{n}\tilde{A}^{+}\bm{1}_{X}$ for each $X\subseteq S$. Lastly, observe that $\tilde{A}^{+}$ has the following properties: $\limsup_{n}\sum_{k}\tilde{a}_{n,k}^{+}=1$, $\lim_{n}\tilde{a}_{n,k}^{+}=0$, and $\sum_{k}\tilde{a}_{n,k}^{+}<\infty$ for all $n,k\in\omega$. In particular, even if $\tilde{A}^{+}$ is not necessarily regular, it satisfies the hypotheses of [8, Theorem 6.2 and Proposition 7.2]. Hence, we proceed verbatim as in Claim 4 and we obtain that $\tilde{\mathcal{I}}$ is not tall. Since $S$ is arbitrary, we conclude that $\mathcal{I}$ is a nowhere tall ideal. ∎
Thanks to Claims 1, 3, and 5 we know that $\mathcal{I}$ is an analytic $P$-ideal which is also nowhere tall (notice that such properties are invariant under isomorphisms). Then, it is known that $\mathcal{I}$ is necessarily $\mathrm{Fin}$ or (an isomorphic copy on $\omega$ of) $\mathrm{Fin}\oplus\mathcal{P}(\omega)$ or {}×$\{\emptyset\}\times\mathrm{Fin}$, see e.g. [9, Corollary 1.2.11] or [11, Theorem 2.26]. Finally, $\mathcal{I}$ has to be also a $P^{+}$-ideal by Claim 2. Hence, it is immediate to check that $\mathcal{I}$ cannot be a copy of {}×$\{\emptyset\}\times\mathrm{Fin}$. This concludes the proof. ∎
Remark 3.4.
Pick an infinite set $T\subseteq\omega$ which is not cofinite and define $\mathcal{J}:=\{S\subseteq\omega:S\cap T\in\mathrm{Fin}\}$ (hence, $\mathcal{J}$ is an isomorphic copy on $\omega$ of $\mathrm{Fin}\oplus\mathcal{P}(\omega)$). Pick also an ideal $\mathcal{I}$ on $\omega$ such that $\mathcal{I}\leq_{\mathrm{RK}}\mathcal{J}$. Thanks to Theorem 2.4, there exists an infinite matrix $A\in(\ell_{\infty},\ell_{\infty})$ such that equality (2) holds. Now, as it has been observed in the proof of the If part of Theorem 2.5, we have also $\mathcal{J}\leq_{\mathrm{RK}}\mathrm{Fin}$. Hence, with the same argument, there exists a matrix $B\in(\ell_{\infty},\ell_{\infty})$ such that $\mathrm{core}_{B\bm{x}}(\mathrm{Fin})=\mathrm{core}_{\bm{x}}(\mathcal{J})$ for all sequences $\bm{x}\in\ell_{\infty}$. In addition, by the proof of Theorem 2.4, it is possible to assume that each row of $B$ contains a single $1$. Set $C:=BA$ (observe that each entry of $C$ is well defined) and note that, if $\bm{x}\in\ell_{\infty}$ then $C\bm{x}=B(A\bm{x})$ is bounded, hence $C\in(\ell_{\infty},\ell_{\infty})$. It follows also that
$\forall\bm{x}\in\ell_{\infty},\quad\mathrm{core}_{C\bm{x}}(\mathrm{Fin})=\mathrm{core}_{A\bm{x}}(\mathcal{J})=\mathrm{core}_{\bm{x}}(\mathcal{I}).$
This does not contradict the claim of Theorem 2.5. In fact, we claim that, if $\mathcal{I}\leq_{\mathrm{RK}}\mathcal{J}$ then either $\mathcal{I}=\mathrm{Fin}$ or $\mathcal{I}$ is an isomorphic copy of $\mathrm{Fin}\oplus\mathcal{P}(\omega)$.
For, let $h:\omega\to\omega$ be a map which witnesses $\mathcal{I}\leq_{\mathrm{RK}}\mathcal{J}$. Observe that $W:=h[T]\in\mathcal{I}^{+}$ since $T\in\mathcal{J}^{+}$, hence $W$ is an infinite set. Considering that $h^{-1}[F]\in\mathrm{Fin}$ if and only if $F\in\mathrm{Fin}$ for each $F\subseteq\omega$, we obtain that
$\begin{split}\forall S\in\omega,\quad S\in\mathcal{I}&\,\,\Longleftrightarrow\,\,h^{-1}[S]\in\mathcal{J}\,\,\Longleftrightarrow\,\,h^{-1}[S]\cap T\in\mathrm{Fin}\\ &\,\,\Longleftrightarrow\,\,h^{-1}[S\cap W]\in\mathrm{Fin}\,\,\Longleftrightarrow\,\,S\cap W\in\mathrm{Fin}.\end{split}$
To sum up, $W$ is an infinite set and =$\mathcal{I}=\{S\subseteq\omega:S\cap W\in\mathrm{Fin}\}$. Therefore either $\mathcal{I}=\mathrm{Fin}$ (in the case where $W$ is cofinite) or $\mathcal{I}$ is isomorphic to $\mathrm{Fin}\oplus\mathcal{P}(\omega)$ (in the opposite case).
4. Acknowledgments
The author is grateful to an anonymous referee for a careful reading of the manuscript and several useful observations.